Please help me with this word problem:

"Find three consecutive integers such that 5 times the largest integer is equal to 22 added to the sum of the first two."
Is it possible i got it right before? this is how i set up the problem the first time:
5(x+2)=22+x+(x+1)


Found 2 solutions by ovie27, Alan3354:
Answer by ovie27(16)

Yes, you are correct.
The largest consecutive integer is x+2
The smallest is x.
5x+10=2x+23
3x=13
x=13/3
The second integer is 13/3 +1 = 16/3
The last one is 19/3.

Answer by Alan3354(69443)

"Find three consecutive integers such that 5 times the largest integer is equal to 22 added to the sum of the first two."
Is it possible i got it right before? this is how i set up the problem the first time:
5(x+2)=22+x+(x+1)
-------------------
Looks right to me.
5x+10 = 22 + 2x + 1 = 2x + 23
3x = 13
There are no such integers.
--> no solution